OpenStudy (anonymous):

Why is Sn(SO4)2 named Tin (IV) Sulfate and not Tin (II) Sulfate?

OpenStudy (anonymous):

Because you have to look at the charges. (SO4)2 has a -5 charge so to be balanced tin has to have a +5 charge. therefore the compound is said to be Tin (IV) Sulfate. This is true for name all compounds with transition metals.

OpenStudy (anonymous):

sorry definitely meant 4. the sulfate ion has a -2 charge so two of them would have a -4 charge so tin has to have a +4 charge.

OpenStudy (abb0t):

charge of SO4 = 2- It might help if you thought of it in terms of a math equation where you're solving for "x". X represents the oxidation state of your metal. You have TWO SO4 which has a formal charge of -2 which mean syou have a total of -4 hence, 2(-2) = -4 the total charge on your atom is zero (it carries no charge). hence it's zero. Remember your molecule total needs to be neutral. Hence +4 - 4 = 0 x + 2(-2) = 0 x - 4 = 0 x = +4 hence, you're oxidation state of Sn = 4